OpenStudy (anonymous):

Which of the following is the solution to the equation square root of 5^8n = 125^(n + 2) ? Choose one answer. a. n = -2 b. n = 2 c. n = -6 d. n = 6

OpenStudy (anonymous):

since \(125=5^3\) you can start with \[5^{8n}=5^{3n+6}\]

OpenStudy (anonymous):

and then go to \[8n=3n+6\] and solve for \(n\)

OpenStudy (anonymous):

I have tried everything i can't do it I can't figure it out

OpenStudy (anonymous):

although i don't see the correct answer in your choices. was the original problem \[5^{8n}=125^{n+2}\]?

OpenStudy (anonymous):

nope its was the sqrt of 5 then ^8n

OpenStudy (anonymous):

oooooooooooooooh ok

OpenStudy (anonymous):

\[\sqrt{5}^{8n}=5^{4n}\]

OpenStudy (anonymous):